The question is labelled W.26. Thank you!

onstruct an analogous five-phase miracle. N.26 C5s frequency is 523 Hz) and both hear beating at a frequency of 3 Hz. Assuming that one of W.27 A piano tuner listens carefully to a tuning fork calibrated to vibrate at 264.13 Hz Two musicians simultaneously play what each hopes is high C (C5s fre pproximately he instruments is precisely in tune, ascertain the possible frequencies of the other.
Answer #1

Frequency of one of the musicians, f1 = 523 Hz

Frequency of the other musician is f2

Beat frequency, delta f = |f2-f1| = 3 Hz

f2-f1 = 3 or f2-f1 = -3

f2 = 526 Hz or f2 = 520 Hz (possible frequencies)